Find three consecutive odd integers such that the product of -7and the sum of the first and third is 27 greater than the product of 11 and the opposite of the second. Please help
Answer by Edwin McCravy(20060)

Least (first) odd integer = x-2
Middle (second) odd integer = x
Greatest (third) odd integer = x+2



Least (first) odd integer = -9-2 = -11
Middle (second) odd integer = -9
Greatest (third) odd integer = -9+2 = -7

Edwin